What is the frequency of green light, which has a wavelength of 4.9X10^-7m?
1
Expert's answer
2020-10-04T09:20:09-0400

The relationship between the wavelength λ and the frequency ν is the following:

"λ = \frac{c}{ ν}"

"ν = \frac{c}{λ}"

c is the speed of light 2.9979 ×108 m/s.

"ν = \frac{2.9979 \times 10^8\;m/s}{4.9 \times 10^{-7}\;m} = 6.11 \times 10^{14}\;1/s = 6.11 \times 10^{14}\;Hz"

Answer: 6.11×1014 Hz